OpenStudy (anonymous):

4(7 - 3x) = 7(4 - 2x)

OpenStudy (lgbasallote):

distribute first \[\implies 28 - 12x = 28 - 14x\] add 14x to both sides \[\implies 28 + 14x - 12x = 28 - 14x + 14x\] combine like terms... \[\implies 28 + 2x = 28\] subtract 28 from both sides \[\implies 2x = 0\] can you solve for x now?

OpenStudy (anonymous):

How do you get all of the fancy signs @lgbasallote

OpenStudy (lgbasallote):

you get some privileges as you level up